0 0
Add a comment Improve this question Transcribed image text
Answer #1

Work Done in the isothermal expansion of an ideal gas is W_i = nRTln[\frac{P_1}{P_2}]

if we model air as an ideal gas, we can write: PV = nRT

therefore, Pi 1V1In P2

P1 = Pgauge + Patm = 110kPa + 101.3kPa = 211.3 kPa

P2 = 101.3kPa

therefore, W_i = [211.3\times 10^3](0.9)ln[\frac{211.3\times 10^3}{101.3\times 10^3}] = 139.81kJ .

P1V1 = P2V2

therefore V2 = (211.3 x 103 x 0.9)/(101.3 x 103) = 1.877 m3

and so the work done in the isobaric process will be: Wf = 101 x 103[0.9 - 1.877] = - 99kJ

therefore the total Work Done will be: W = 139.81 + ( - 99) = 40.81 kJ.